b) The current (Year 2017) atmospheric CO2 concentration is 370 ppm. It is estimated that if we continue to exploit the world’s fossil fuel reserves at the same rate, the atmospheric CO2 concentration will reach 1,100 ppm in Year 2400 (Williams, 2001).

Justify your answer in (1) quantitatively by calculating the pHs when CO2 concentrations are 370 ppm (Year 2017) and 1,100 ppm (Year 2400), respectively. Assuming: I. Pure water is in equilibrium with CO2 at 25oC.
II. Water vapor pressure is 0.031 atm.
III. Henry’s Law constant 3.38x10-2 (mol/L) x atm-1 at 25oC for CO2.
IV. Ka1 = 4.45x10-7; Ka2 = 4.69x10-11
Neglect all other processes such as complexation, and precipitation.
Answer #1

H2O + CO2 <----> H2CO3(aq)

According to the ideal gas equation: PV = nRT, then P = CRT

Where 'C' is the concentration of CO2 = 370 ppm = 370 mg/L = 0.00841 mol/L

Note: Molar mass of CO2, i.e. P(CO2) = 44 g/mol, i.e. moles of CO2 = 370 mg/44 g.mol-1 = 8.41 mol or 0.00841 mmol

Now, pressure of CO2 = 0.00841 mol/L * 0.0821 L.atm/mol.K * (25+273) K = 0.20574 atm

According to the Henry's law:

The concentration of the gas in water = KH * P(CO2) = 3.38*10-2 mol/L.atm * 0.20574 atm = 0.006954 mol/L or 0.006954 M